# Character theoretic techniques for nonabelian partial difference sets

### v1

@@ Abstract.
A $(v,k,\lambda,\mu)$-partial difference set (PDS) is a subset $D$ of size $k$ of a group $G$ of order $v$ such that every nonidentity element $g$ of $G$ can be expressed in either $\lambda$ or $\mu$ different ways as a product $xy^{-1}$, $x,y\in D$, depending on whether or not $g$ is in $D$. If $D$ is inverse closed and $1\notin D$, then the Cayley graph ${\rm Cay}(G,D)$ is a $(v,k,\lambda,\mu)$-strongly regular graph (SRG). PDSs have been studied extensively over the years, especially in abelian groups, where techniques from character theory have proven to be particularly effective. Recently, there has been considerable interest in studying PDSs in nonabelian groups, and the purpose of this paper is develop character theoretic techniques that apply in the nonabelian setting. We prove that analogues of character theoretic results of Ott [25] about generalized quadrangles of order $s$ also hold in the general PDS setting, and we are able to use these techniques to compute the intersection of a putative PDS with the conjugacy classes of the parent group in many instances. With these techniques, we are able to prove the nonexistence of PDSs in numerous instances and provide severe restrictions in cases when such PDSs may still exist. Furthermore, we are able to use these techniques constructively, computing several examples of PDSs in nonabelian groups not previously recognized in the literature, including an infinite family of genuinely nonabelian PDSs associated to the block-regular Steiner triple systems originally studied by Clapham [6] and related infinite families of genuinely nonabelian PDSs associated to the block-regular Steiner 222-designs first studied by Wilson [33].
@@ 1. Introduction
If $G$ is a group, then we say that $D\subset G$ is a $(v,k,\lambda,\mu)$-partial difference set (PDS) if $G$ has order $v$, $D$ has size $k$, and if the product $gh^{-1}$ for all $g,h\in D$ with $g\neq h$ represents all nonidentity elements in $D$ exactly $\lambda$ times and all nonidentity elements in $G-D$ exactly $\mu$ times. We refer to the tuple $(v,k,\lambda,\mu)$ as the parameters of $D$. In addition, we call a PDS regular if $D$ is closed under taking inverses (in which case we write $D=D^{(-1)}$) and $1\not\in D$. This definition is not particularly restrictive, for if $D$ is a PDS with $1\in D$, then $D-\{1\}$ is a PDS as well, and if $\lambda\neq\mu$, then $D=D^{(-1)}$. If we have $\lambda=\mu$ then $D$ is called a difference set, and we say that $D$ is a $(v,k,\lambda)$-difference set (DS). We additionally call $D$ a reversible $(v,k,\lambda)$-DS if $D=D^{(-1)}$.
@@
[…]
@@
Recently, Ott [25] studied $\Phi$ – which is indeed a class function on $G$ – from a character-theoretic perspective, and Ott used this to prove that GQs of order $s$ ($s+1$ points incident with each line, $s+1$ lines incident with each point), $s$ even, cannot have a group acting regularly on its point set. As it turns out, the ideas Ott used are applicable in situations far beyond GQs (and reversible difference sets); in fact, one could argue that the PDS setting is in fact where these techniques are most effective.
@@
We have used these methods to show that if a $(v,k,\lambda,\mu)$-PDS exists in a group $G$, where there is a minimal nontrivial normal subgroup $N$ of $G$ such that $G/N$ is abelian and $\gcd(|G/N|,\sqrt{\Delta})=1$, then the size of the intersection of the PDS with the cosets of $N$ are known, providing severe restrictions on such PDSs. We have also used our methods to construct several PDSs. For $p$ a prime, we demonstrate the existence of an infinite family of PDSs in nonabelian groups related to Steiner triple systems with parameters $(p^{d}(p^{d}-1)/6,3(p^{d}-3)/2,(p^{d}+3)/2,9)$, where $p^{d}$ is greater than 999 and congruent to $7\pmod{12}$, as well as Steiner 222-designs ${\mathrm{S}}(2,k,v)$ with parameters $(p^{d}(p^{d}-1)/(k(k-1)),k(p^{d}-k)/(k-1),(p^{d}-1)/(k-1)+(k-1)^{2}-2,k^{2})$ for $p^{d}$ sufficiently large and $p^{d}\equiv k(k-1)+1\pmod{2k(k-1)}$. (These Steiner triple systems were shown to be block-regular by Clapham [6], and the Steiner 222-designs were shown to be block regular by Wilson [33]. It is immediate that these will be PDSs; however, PDS families in Steiner 222-designs do not seem to have been recognized thus far in the literature beyond a recent mention in [28].)
This paper is organized as follows. In Section 2, we outline preliminary results related to PDSs, SRGs, and character theory over local rings necessary for later sections. Section 3 is dedicated to generalizing the results of Ott to the PDS setting. In Section 4, we introduce computational methods for the case when one of $\sqrt{\Delta}$, $v$ has factors which are relatively prime to the others that are used both to rule out the existence of numerous PDSs and also discover several new examples in nonabelian groups; we also provide an example demonstrating how the techniques from this paper can be used when $\sqrt{\Delta}$ and $v$ have the same primes as factors (such as when $\Delta=v$). Finally, we include appendices in which we list parameters ruled out either theoretically or computationally (Appendices A, B) and parameters for which there exists a potentially valid $\Phi$ function (Appendix C).
@@ 2.2. Strongly regular graphs
[…]
@@
Let $\Gamma$ be a $(v,k,\lambda,\mu)$-SRG with eigenvalues $k,\theta_{1},\theta_{2}$ as before. Then,
@@
[…]
@@ 2.3. Some families of SRGs
[…]
One particularly important family of strongly regular graphs are those related to Steiner 222-designs, which we use in Proposition 4.9, Theorem 4.11, Proposition 4.12, and Proposition 4.13. A Steiner 222-design, ${\mathrm{S}}(2,k,v)$, is a $v$-element set, along with a collection of $k$-subsets of this set, called blocks, such that any two elements are contained in exactly one block. The families of Steiner 222-designs in Proposition 4.9, Theorem 4.11, Propostion 4.12, and Proposition 4.13 are all algebraic constructions using the elements of an abelian group $G$. In these cases, the Steiner 222-design is represented by a distinguished set $\mathscr{B}$ of blocks, known as base blocks. The remaining blocks take the form $g+B$ for $B\in\mathscr{B}$ and $g\in G$.
A Steiner 222-design has a natural representation as a strongly regular graph if one identifies the blocks of the Steiner 222-design with the vertices of the graph such that two vertices are adjacent if the intersection of their corresponding blocks has one element. If $v$ is the number of elements in the Steiner 222-design, then the corresponding SRG has parameters
@@
We will speak often of one particular class of Steiner 222-designs, known as Steiner triple systems, which occur when $k=3$. For further information on Steiner triple systems, Steiner 222-designs, and Steiner systems more generally, see [7, Section I.4].
@@ 2.4. Character theory and representations over local rings
=$\mathfrak{R}=\biggl{\{}\frac{x}{y}:x,y\in\mathbb{Z}[\zeta],y\not\in\mathfrak{p}_{0}\biggr{\}}.$
@@ 3.1. Results that hold for any $v$, $\sqrt{\Delta}$
For $D$ a regular PDS inside $G$ and a representation $\mathfrak{F}$ affording a nonprincipal irreducible character $\chi$ with underlying vector space $V_{\mathfrak{F}}$, $\mathfrak{F}(D)$ has two eigenvalues $\theta_{1}=\frac{1}{2}(\lambda-\mu+\sqrt{\Delta})$ and $\theta_{2}=\frac{1}{2}(\lambda-\mu-\sqrt{\Delta})$ with corresponding eigenspaces $V_{\mathfrak{F}}(\theta_{1})$ and $V_{\mathfrak{F}}(\theta_{2})$ such that $V_{\mathfrak{F}}=V_{\mathfrak{F}}(\theta_{1})\oplus V_{\mathfrak{F}}(\theta_{2})$, and
@@
0 0 $=\mathfrak{F}(D^{2})-(k-\mu)I-(\lambda-\mu)\mathfrak{F}(D).$
@@
0 0 $=\Bigl{(}\mathfrak{F}(D)^{2}-(\lambda-\mu)\mathfrak{F}(D))-(k-\mu)I\Bigr{)}v$
0 0 $=\alpha^{2}+(\mu-\lambda)\alpha+(\mu-k).$
@@
We have just shown that $\mathfrak{F}(D)$ is a solution to the polynomial $z^{2}-(\lambda-\mu)z+\mu-k$, where $z^{2}-(\lambda-\mu)z+\mu-k=(z-\theta_{1})(z-\theta_{2})$ by Lemma 2.2. Moreover, $\theta_{1}\neq\theta_{2}$. For if $\theta_{1}=\theta_{2}$, then we must have $\mu\geqslant k$, but this is impossible, for $\mu\leqslant k$ by definition, and if $\mu=k$, then $\lambda=\mu$ cannot hold. So, $\theta_{1}\neq\theta_{2}$. Therefore, the minimal polynomial of $\mathfrak{F}(D)$ is degree 111 or 222 with distinct roots, so $\mathfrak{F}(D)$ is diagonalizable with $\theta_{1},\theta_{2}$ on the diagonal. If $V_{\mathfrak{F}}(\theta_{i})$ is the eigenspace corresponding to the eigenvectors of $\theta_{i}$, then $V_{\mathfrak{F}}=V_{\mathfrak{F}}(\theta_{1})\oplus V_{\mathfrak{F}}(\theta_{2})$ and $\chi(D)=\dim V_{\mathfrak{F}}(\theta_{1})\theta_{1}+\dim V_{\mathfrak{F}}(\theta_{2})\theta_{2}$. Note that if $\chi$ is a linear character, then it has one eigenspace of dimension 111, in which case $\chi(D)=\theta_{1}$ or $\chi(D)=\theta_{2}$. ∎
@@
Assume ${\rm Cay}(G,D)$ is not a conference graph and $\mu>0$. Let $\xi\in\mathcal{L}$ be a nonprincipal linear character of order $p$ with kernel $N$. Then, $p|(k-\xi(D))$, $|N\cap D|=\frac{k-\xi(D)}{p}+\xi(D)$, and $|Na\cap D|=\frac{k-\xi(D)}{p}$ for every $a\not\in N$.
@@
[…]
@@
[…]
@@ 3.2. Results that hold when a prime $p$ divides $v$ but not $\sqrt{\Delta}$
Suppose $p$ is a prime dividing $v$ but not $\sqrt{\Delta}$. Assuming $\mu>0$, there exists a factorization $\mu=\mu_{1}\mu_{2}$ such that $v_{1}=(k-\theta_{1})/\mu_{1}$, $v_{2}=(k-\theta_{2})/\mu_{2}$, $v=v_{1}v_{2}$, and $p$ divides only one of $v_{1}$, $v_{2}$.
@@
Let $G$ be a group with a regular PDS $D$ and a nontrivial group of linear characters $\mathcal{L}$. Let $P\subseteq\mathcal{L}$ be a Sylow $p$-subgroup of $\mathcal{L}$, and suppose further that $p$ and $\sqrt{\Delta}$ are coprime. Then, for any $\xi^{\prime}\in P$,
@@
[…]
@@
[…]
@@
[…]
@@
Let $\xi\in\mathcal{L}$ be a linear character of order $p^{a}$ with $p$ coprime to $\sqrt{\Delta}$, and let $\chi\in{\rm Irr}(G)$. Then, $\xi\chi(D)=\chi(D)$.
The following proof is quite technical, so we provide a broad explanation of its various moving parts. We would like to show that $\chi(D)=\xi\chi(D)$. To do this, we use Theorem 2.5 to take a $\mathbb{Q}(\zeta)$-representation $\mathfrak{F}$ affording the $\mathbb{C}$-character $\chi$, use Lemma 2.4 to assume that $\mathfrak{F}$ is a $\mathfrak{R}$-representation, and construct a homomorphism, which we denote by $\overline{\phantom{x}}$, from the $\mathfrak{R}$-module underlying $\mathfrak{F}$ to an $L$-module, which is also an $L$-vector space since $L$ is a field. Under this homomorphism, $\overline{\xi(D)\mathfrak{F}(D)}=\overline{\mathfrak{F}(D)}$, since for all $x\in G$, $\overline{\xi(x)}$ is a scalar whose order divides $p^{a}$ in $L$, so $\overline{\xi(x)}$ must evaluate to 111. We then show that for any matrix $\mathfrak{F}(D)$ with $\mathbb{Q}(\zeta)$-eigenspace $V_{\mathfrak{F}}(\theta_{\alpha})$ and $\mathfrak{R}$-eigenspace $M_{\mathfrak{F}}(\theta_{\alpha})$, that
@@
[…]

### v2

@@ Abstract.
A $(v,k,\lambda,\mu)$-partial difference set (PDS) is a subset $D$ of size $k$ of a group $G$ of order $v$ such that every nonidentity element $g$ of $G$ can be expressed in either $\lambda$ or $\mu$ different ways as a product $xy^{-1}$, $x,y\in D$, depending on whether or not $g$ is in $D$. If $D$ is inverse closed and $1\notin D$, then the Cayley graph ${\rm Cay}(G,D)$ is a $(v,k,\lambda,\mu)$-strongly regular graph (SRG). PDSs have been studied extensively over the years, especially in abelian groups, where techniques from character theory have proven to be particularly effective. Recently, there has been considerable interest in studying PDSs in nonabelian groups, and the purpose of this paper is develop character theoretic techniques that apply in the nonabelian setting. We prove that analogues of character theoretic results of Ott [25] about generalized quadrangles of order $s$ also hold in the general PDS setting, and we are able to use these techniques to compute the intersection of a putative PDS with the conjugacy classes of the parent group in many instances. With these techniques, we are able to prove the nonexistence of PDSs in numerous instances and provide severe restrictions in cases when such PDSs may still exist. Furthermore, we are able to use these techniques constructively, computing several examples of PDSs in nonabelian groups not previously recognized in the literature, including an infinite family of genuinely nonabelian PDSs associated to the block-regular Steiner triple systems originally studied by Clapham [6] and related infinite families of genuinely nonabelian PDSs associated to the block-regular Steiner $2$-designs first studied by Wilson [33].
@@ 1. Introduction
If $G$ is a group, then we say that $D\subset G$ is a $(v,k,\lambda,\mu)$-partial difference set (PDS) if $G$ has order $v$, $D$ has size $k$, and if the product $gh^{-1}$ for all $g,h\in D$ with $g\neq h$ represents all nonidentity elements in $D$ exactly $\lambda$ times and all nonidentity elements in GDG-D exactly $\mu$ times. We refer to the tuple $(v,k,\lambda,\mu)$ as the parameters of $D$. In addition, we call a PDS regular if $D$ is closed under taking inverses (in which case we write $D=D^{(-1)}$) and $1\not\in D$. This definition is not particularly restrictive, for if $D$ is a PDS with $1\in D$, then $D-\{1\}$ is a PDS as well, and if $\lambda\neq\mu$, then $D=D^{(-1)}$. If we have $\lambda=\mu$ then $D$ is called a difference set, and we say that $D$ is a $(v,k,\lambda)$-difference set (DS). We additionally call $D$ a reversible $(v,k,\lambda)$-DS if $D=D^{(-1)}$.
@@
where $\Delta\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}(\lambda-\mu)^{2}+4(k-\mu)$ (see [24, Corollary 3.3]).
@@
Recently, Ott [25] studied $\Phi$ – which is indeed a class function on $G$ – from a character-theoretic perspective, and Ott used this to prove that GQs of order $s$ ($s+1$ points incident with each line, $s+1$ lines incident with each point), $s$ even, cannot have a group acting regularly on its point set. As it turns out, the ideas Ott used are applicable in situations far beyond GQs (and reversible difference sets); in fact, one could argue that the PDS setting is where these techniques are most effective.
@@
We have used these methods to show that if a $(v,k,\lambda,\mu)$-PDS exists in a group $G$, where there is a minimal nontrivial normal subgroup $N$ of $G$ such that $G/N$ is abelian and $\gcd(|G/N|,\sqrt{\Delta})=1$, then the size of the intersection of the PDS with the cosets of $N$ are known, providing severe restrictions on such PDSs. We have also used our methods to construct several PDSs. For $p$ a prime, we demonstrate the existence of an infinite family of PDSs in nonabelian groups related to Steiner triple systems with parameters $(p^{d}(p^{d}-1)/6,3(p^{d}-3)/2,(p^{d}+3)/2,9)$, where $p^{d}$ is greater than $9$ and congruent to $7\pmod{12}$, as well as Steiner $2$-designs ${\mathrm{S}}(2,k,v)$ with parameters $(p^{d}(p^{d}-1)/(k(k-1)),k(p^{d}-k)/(k-1),(p^{d}-1)/(k-1)+(k-1)^{2}-2,k^{2})$ for $p^{d}$ sufficiently large and $p^{d}\equiv k(k-1)+1\pmod{2k(k-1)}$. (These Steiner triple systems were shown to be block-regular by Clapham [6], and the Steiner $2$-designs were shown to be block regular by Wilson [33]. It is immediate that these will be PDSs; however, PDS families in Steiner $2$-designs do not seem to have been recognized thus far in the literature beyond a recent mention in [28].)
This paper is organized as follows. In Section 2, we outline preliminary results related to PDSs, SRGs, and character theory over local rings necessary for later sections. Section 3 is dedicated to generalizing the results of Ott to the PDS setting. In Section 4, we introduce computational methods for the case when one of $\sqrt{\Delta}$, $v$ has factors which are relatively prime to the others that are used both to rule out the existence of numerous PDSs and also discover several new examples in nonabelian groups; we also provide an example demonstrating how the techniques from this paper can be used when $\sqrt{\Delta}$ and $v$ have the same primes as factors (such as when $\Delta=v$). Finally, we include appendices in which we list parameters that can be ruled out either theoretically or computationally by methods known already known (Appendix A) or by new methods (Appendix B) as well as a list of open cases and cases in which a PDS exists for sufficiently small $v$ (Appendix C).
@@ 2.2. Strongly regular graphs
where $\Delta\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}(\lambda-\mu)^{2}+4(k-\mu)$; we call this value the discriminant of the PDS. The eigenvalue $k$ is known as the Perron eigenvalue, and it has multiplicity $1$. The eigenvalues $\theta_{1}$ and $\theta_{2}$ occur with multiplicities $m_{1}$ and $m_{2}$ in the adjacency matrix of the associated SRG, where
@@
Let $\Gamma$ be a $(v,k,\lambda,\mu)$-SRG with eigenvalues $k,\theta_{1},\theta_{2}$. Then,
@@
Note that for our purposes $\sqrt{\Delta}$ will always be an integer. Indeed, the case when $\sqrt{\Delta}$ is not an integer is quite restricted: the SRG $\Gamma$ must be a conference graph, which has parameters (v,(v1)/2,(v5)/4,(v1)/4)(v,(v-1)/2,(v-5)/4,(v-1)/4); see [16, Lemmas 10.3.2, 10.3.3].
@@ 2.3. Some families of SRGs
We denote a $\rm GQ$} of order $(s,t)$ – which may not be unique – by ${\rm GQ}(s,t)$. We may construct an incidence graph $\Gamma$ from a GQ $S$ by identifying the points of $S$ with the vertices of $\Gamma$, and specifying that two points are adjacent if they are collinear. It is an easy exercise that $\Gamma$ is a $($s+1)(st+1),s(t+1),s1,t+1)$($s+1)(st+1),s(t+1),s-1,t+1)-SRG. See [4, Section 2.2.9] for further information on GQs.
One particularly important family of strongly regular graphs are those related to Steiner $2$-designs, which we use in Proposition 4.9, Theorem 4.11, Proposition 4.12, and Proposition 4.13. A Steiner $2$-design, ${\mathrm{S}}(2,k,v)$, is a $v$-element set, along with a collection of $k$-subsets of this set, called blocks, such that any two elements are contained in exactly one block. The families of Steiner $2$-designs in Proposition 4.9, Theorem 4.11, Propostion 4.12, and Proposition 4.13 are all algebraic constructions using the elements of an abelian group $G$. In these cases, the Steiner $2$-design is represented by a distinguished set $\mathscr{B}$ of blocks, known as base blocks. The remaining blocks take the form $g+B$ for $B\in\mathscr{B}$ and $g\in G$.
A Steiner $2$-design has a natural representation as a strongly regular graph if one identifies the blocks of the Steiner $2$-design with the vertices of the graph such that two vertices are adjacent if the intersection of their corresponding blocks has one element. If $v$ is the number of elements in the Steiner $2$-design, then the corresponding SRG has parameters
@@
We will speak often of one particular class of Steiner $2$-designs, known as Steiner triple systems, which occur when $k=3$. For further information on Steiner triple systems, Steiner $2$-designs, and Steiner systems more generally, see [7, Section I.4].
@@ 2.4. Character theory and representations over local rings
$\mathfrak{R}=\biggl\{\frac{x}{y}:x,y\in\mathbb{Z}[\zeta],y\not\in\mathfrak{p}_{0}\biggr\}.$
@@ 3.1. Results that hold for any $v$, $\sqrt{\Delta}$
For $D$ a regular $(v,k,\lambda,\mu)$-PDS inside $G$ and a representation $\mathfrak{F}$ affording a nonprincipal irreducible character $\chi$ with underlying vector space $V_{\mathfrak{F}}$, $\mathfrak{F}(D)$ has two eigenvalues $\theta_{1}=\frac{1}{2}(\lambda-\mu+\sqrt{\Delta})$ and $\theta_{2}=\frac{1}{2}(\lambda-\mu-\sqrt{\Delta})$ with corresponding eigenspaces $V_{\mathfrak{F}}(\theta_{1})$ and $V_{\mathfrak{F}}(\theta_{2})$ such that $V_{\mathfrak{F}}=V_{\mathfrak{F}}(\theta_{1})\oplus V_{\mathfrak{F}}(\theta_{2})$, and
@@
$0$ $=\mathfrak{F}(D^{2})-(k-\mu)I-(\lambda-\mu)\mathfrak{F}(D).$
@@
$0$ $=\Bigl(\mathfrak{F}(D)^{2}-(\lambda-\mu)\mathfrak{F}(D))-(k-\mu)I\Bigr)v$
$0$ $=\alpha^{2}+(\mu-\lambda)\alpha+(\mu-k).$
@@
We have just shown that $\mathfrak{F}(D)$ is a solution to the polynomial $z^{2}-(\lambda-\mu)z+\mu-k$, where $z^{2}-(\lambda-\mu)z+\mu-k=(z-\theta_{1})(z-\theta_{2})$ by Lemma 2.2. Moreover, $\theta_{1}\neq\theta_{2}$. For if $\theta_{1}=\theta_{2}$, then we must have $\mu\geqslant k$, but this is impossible, for $\mu\leqslant k$ by definition, and if $\mu=k$, then $\lambda=\mu$ cannot hold. So, $\theta_{1}\neq\theta_{2}$. Therefore, the minimal polynomial of $\mathfrak{F}(D)$ is degree $1$ or $2$ with distinct roots, so $\mathfrak{F}(D)$ is diagonalizable with $\theta_{1},\theta_{2}$ on the diagonal. If $V_{\mathfrak{F}}(\theta_{i})$ is the eigenspace corresponding to the eigenvectors of $\theta_{i}$, then $V_{\mathfrak{F}}=V_{\mathfrak{F}}(\theta_{1})\oplus V_{\mathfrak{F}}(\theta_{2})$ and $\chi(D)=\dim V_{\mathfrak{F}}(\theta_{1})\theta_{1}+\dim V_{\mathfrak{F}}(\theta_{2})\theta_{2}$. Note that if $\chi$ is a linear character, then it has one eigenspace of dimension $1$, in which case $\chi(D)=\theta_{1}$ or $\chi(D)=\theta_{2}$. ∎
@@
Assume $G$ contains a regular $(v,k,\lambda,\mu)$-PDS with nonprincipal eigenvalues $\theta_{1}$, $\theta_{2}$, ${\rm Cay}(G,D)$ is not a conference graph, and $\mu>0$. Let $\mathcal{L}$ be a nontrivial group of linear characters of $G$, and let $\xi\in\mathcal{L}$ be a nonprincipal linear character of order $p$ with kernel $N$. Then, $p|(k-\xi(D))$, $|N\cap D|=\frac{k-\xi(D)}{p}+\xi(D)$, and $|Na\cap D|=\frac{k-\xi(D)}{p}$ for every $a\not\in N$.
@@
    We again remark that the results in this subsection do not require that $G$ be an abelian or nonabelian group or any conditions on the factors of $|G|$ or $\sqrt{\Delta}$. Thus, these results may be applied even in cases when our techniques in the rest of the paper do not otherwise work; an example of such an application is given in Subsection 4.4
@@
    The main technique in [25] and our own Theorem 3.8 is to extract information about $D$ by analyzing $\mathfrak{F}(D)$ over fields of characteristic $p$, when $p$ does not divide $\sqrt{\Delta}=\theta_{1}-\theta_{2}$ but does divide $|G|$. A critical element in this proof is that $\mathfrak{F}(D)$ induces a representation on a certain field $L$ of characteristic $p$, and that the induced representation $\overline{\mathfrak{F}(D)}$ on $L$ is diagonalizable, for its minimal polynomial is $(z-\theta_{1})(z-\theta_{2})\pmod{p}$, where $\theta_{1}\not\equiv\theta_{2}\pmod{p}$. As a result, while the results of this subsection still hold, the techniques in the following subsections largely fail if we ever have $|G|$ sharing all of its prime divisors with $\theta_{1}-\theta_{2}$, for then the minimal polynomial of $\mathfrak{F}(D)$ for any representation $\mathfrak{F}$ is $(z-\theta_{1})(z-\theta_{2})\equiv(z-\theta_{1})^{2}\pmod{p}$, so $\mathfrak{F}(D)$ is not diagonalizable. Unfortunately, this case happens reasonably often, so for such parameter sets techniques involving local rings are ineffective. In Subsection 4.4, we provide an example computation for how character theory may still be applied to PDSs where $\sqrt{\Delta}$ does divide $v$.
@@ 3.2. Results that hold when a prime $p$ divides $v$ but not $\sqrt{\Delta}$
Let $G$ be a finite group containing a regular $(v,k,\lambda,\mu)$-PDS $D$ that has nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, and let $\sqrt{\Delta}=\theta_{1}-\theta_{2}$. Suppose $p$ is a prime dividing $v$ but not $\sqrt{\Delta}$. Assuming $\mu>0$, there exists a factorization $\mu=\mu_{1}\mu_{2}$ such that $v_{1}=(k-\theta_{1})/\mu_{1}$, $_{2}=(k-\theta_{2})/\mu_{2}, v=v$$v_{1}v_{2}$, and $p$ divides only one of $v_{1}$, $v_{2}$.
@@
Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, and let $\mathcal{L}$ be the group of linear characters of $G$. Let $P\subseteq\mathcal{L}$ be a Sylow $p$-subgroup of $\mathcal{L}$, and suppose further that $p$ and $\sqrt{\Delta}=\theta_{1}-\theta_{2}$ are coprime. Then, for any nontrivial $\xi^{\prime}\in P$,
@@
First, $|P|$ divides $k-\xi(D)$ by Lemmas 3.6 and 3.3. Set $S\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\sum_{\xi\in P}\xi(D)$. Then,
@@
Fix $\theta_{\alpha}$ to be an output of at least one $\xi\in P$ on $D$, and let $r$ of the remaining $\xi$ evaluate to $\theta_{\beta}$ on $D$. Then there are |P|r1|P|-r-1 eigenvalues $\theta_{\alpha}$ in the sum. So,
@@
Recall that $|P|$ divides $k-\xi(D)$. Thus, evaluating $k-\theta_{\alpha}+r\theta_{\beta}+(|P|-r)\theta_{\alpha}$ modulo $|P|$ gives $0\equiv r\theta_{\beta}-r\theta_{\alpha}\equiv\pm r\sqrt{\Delta}$. Further, $\sqrt{\Delta}$ and $|P|$ are coprime, so $r=|P|$ or $r=0$. If $r=|P|$, then every $\xi$ evaluates to $\theta_{\beta}$, which contradicts our selection of $\theta_{\alpha}$ to be the output of some $\xi$, so $r=0$. Therefore, all the linear characters of $P$ are equivalent over $D$. ∎
@@
Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, and let $\mathcal{L}$ be the group of linear characters of $G$. Let $\xi\in\mathcal{L}$ be a linear character of order $p^{a}$ with $p$ coprime to $\sqrt{\Delta}=\theta_{1}-\theta_{2}$, and let $\chi\in{\rm Irr}(G)$. Then, $\xi\chi(D)=\chi(D)$.
The following proof is quite technical, so we provide a broad explanation of its various moving parts. We would like to show that $\chi(D)=\xi\chi(D)$. To do this, we use Theorem 2.5 to take a $\mathbb{Q}(\zeta)$-representation $\mathfrak{F}$ affording the $\mathbb{C}$-character $\chi$, use Lemma 2.4 to assume that $\mathfrak{F}$ is a $\mathfrak{R}$-representation, and construct a homomorphism, which we denote by $\overline{\phantom{x}}$, from the $\mathfrak{R}$-module underlying $\mathfrak{F}$ to an $L$-module, which is also an $L$-vector space since $L$ is a field. Under this homomorphism, $\overline{\xi(D)\mathfrak{F}(D)}=\overline{\mathfrak{F}(D)}$, since for all $x\in G$, $\overline{\xi(x)}$ is a scalar whose order divides $p^{a}$ in $L$, so $\overline{\xi(x)}$ must evaluate to $1$. We then show that for any matrix $\mathfrak{F}(D)$ with $\mathbb{Q}(\zeta)$-eigenspace $V_{\mathfrak{F}}(\theta_{\alpha})$ and $\mathfrak{R}$-eigenspace $M_{\mathfrak{F}}(\theta_{\alpha})$, that
@@
We define the submodule $\mathfrak{p}M_{\mathfrak{F}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\{w\in M_{\mathfrak{F}}:w=qv,q\in\mathfrak{p},v\in M_{\mathfrak{F}}\}$. This is a closed subspace of $M_{\mathfrak{F}}$, so $M_{\mathfrak{F}}/\mathfrak{p}M_{\mathfrak{F}}$ is well-defined. We therefore define the natural projection $\overline{\phantom{x}}:M_{\mathfrak{F}}\longrightarrow M_{\mathfrak{F}}/\mathfrak{p}M_{\mathfrak{F}}$. For an $\mathfrak{R}$-representation $\mathfrak{F}$, there is a corresponding representation $\overline{\mathfrak{F}}$ on $M_{\mathfrak{F}}/\mathfrak{p}M_{\mathfrak{F}}$ by $\overline{\mathfrak{F}(g)}\overline{w}=\overline{\mathfrak{F}(g)w}$. As before, $\overline{\xi(g)}\equiv 1\pmod{\mathfrak{p}}$, so for any vector $w\in M_{\mathfrak{F}}$, we have $\overline{\xi(g)}\overline{w}=\overline{w}$. Thus, for all $w\in V_{\mathfrak{F}}$,